OpenStudy (anonymous):

you are reducing the fraction 80/144 and you mistakenly think the factor you reduce by is four. How can you still get a correctly reduced fraction?

OpenStudy (anonymous):

if you get the greatest common factor of the two numbers you can find the prime factors: 80 = 2*2*2*2*5 144 = 2*2*2*2*3*3 now pair up the factors gcf = 2*2*2*2 = 16 whats left?? 5 and 3*3 : 5 and 9 redused fraction = 5/9